# Mixed precision sketching for least-squares problems and its application in GMRES-based iterative refinement

### v1

@@ 5 Numerics for dense problems
Table 1: Condition numbers and norms when precisions $u_{s}$ and $u_{QR}$ in Algorithm 3.1 are set to half and single, respectively.
$\kappa_{2}(A)$ $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ $\|A\widehat{R}^{-1}\|_{2}$ $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\kappa_{2}(A\widehat{R}^{-1})$
1e+00 to 1e+06 7.45e+00 2.03e+00 1.46e+00 2.97e+00
1e+07 7.65e+00 2.03e+00 1.49e+00 3.01e+00
Table 2: As in Table 1, but both $u_{s}$ and $u_{QR}$ are set to single. $\kappa_{2}(\widetilde{A})$ is as in Table 1.
$\kappa_{2}(A)$ $\kappa_{2}(\widetilde{A})$ $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ $\|A\widehat{R}^{-1}\|_{2}$ $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\kappa_{2}(A\widehat{R}^{-1})$
1e+02 1.62e+04 7.45e+00 2.02e+00 1.46e+00 2.96e+00
1e+04 1.62e+08 3.40e+02 1.80e+00 1.19e+01 2.15e+01
1e+06 1.62e+12 1.81e+06 1.67e+00 8.98e+02 1.50e+03
1e+08 1.39e+16 1.58e+10 1.65e+00 8.43e+04 1.39e+05
1e+10 7.00e+17 1.31e+14 1.55e+00 7.82e+06 1.22e+07
1e+12 1.11e+17 4.81e+17 1.56e+00 6.98e+08 1.09e+09
1e+14 1.44e+17 1.03e+17 1.47e+00 6.32e+10 9.30e+10
1e+15 1.75e+17 1.30e+17 1.49e+00 6.72e+11 1.00e+12
Table 3: As in Table 1, but $u_{s}$ is set to half and $u_{QR}$ is set to double.
$\kappa_{2}(A)$ $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ $\|A\widehat{R}^{-1}\|_{2}$ $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\kappa_{2}(A\widehat{R}^{-1})$
1e+02 to 1e+06 7.45e+00 2.03e+00 1.46e+00 2.97e+00
1e+08 5.64e+01 1.90e+00 4.67e+00 8.92e+00
1e+10 3.98e+05 1.81e+00 4.09e+02 7.41e+02
1e+12 3.37e+09 1.80e+00 3.77e+04 6.80e+04
1e+14 2.71e+13 1.69e+00 3.46e+06 5.86e+06
1e+15 2.71e+15 1.67e+00 3.51e+07 5.86e+07
Table 4: As in Table 1, but $u_{s}$ is set to single and $u_{QR}$ is set to double. $\kappa_{2}(\widetilde{A})$ is as in Table 3.
$\kappa_{2}(A)$ $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ $\|A\widehat{R}^{-1}\|_{2}$ $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\kappa_{2}(A\widehat{R}^{-1})$
1e+02 to 1e+14 7.45e+00 2.03e+00 1.46e+00 2.97e+00
1e+15 7.43e+00 2.03e+00 1.46e+00 2.96e+00
Table 5: As in Table 1, but both $u_{s}$ and $u_{QR}$ are set to double. $\kappa_{2}(\widetilde{A})$ is as in Table 3.
@@
Table 8: LSIR iterations and the total count of FGMRES iterations within LSIR (Algorithm 4.1). Here $u$ is set to single. - denotes that LSIR did not converge in 30 iterations. * denotes when $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES are set to double.
@@
Table 9: As in Table 8, but $u$ is set to double. - denotes that LSIR did not converge in 30 iterations. * denotes when $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES are set to quad and the maximum number of FGMRES iterations is increased to 80.

### v2

@@ 5 Numerics for dense problems
Figure 3: Condition numbers and norms when precisions $u_{s}$ and $u_{QR}$ in Algorithm 3.1 are set to half/single and single, respectively. $\kappa_{2}(\widetilde{A})$ is the same in both cases. Note the different scales for the y axis.
$10^{1}$$10^{4}$$10^{7}$$10^{10}$$10^{13}$$10^{16}$$10^{0}$$10^{3}$$10^{6}$$10^{9}$$10^{12}$$10^{15}$$10^{18}$$\kappa_{2}(A)$$\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$$\|A\widehat{R}^{-1}\|_{2}$$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$$\kappa_{2}(A\widehat{R}^{-1})$$\kappa_{2}(\widetilde{A})$
(a) $u_{s}$ half, $u_{QR}$ double
$10^{1}$$10^{4}$$10^{7}$$10^{10}$$10^{13}$$10^{16}$$10^{0}$$10^{3}$$10^{6}$$10^{9}$$10^{12}$$10^{15}$$10^{18}$$\kappa_{2}(A)$
(b) $u_{s}$ single, $u_{QR}$ double
[…]
(c) $u_{s}$ double, $u_{QR}$ double
Figure 4: As in Figure 3, but $u_{QR}$ is set to double, and $u_{s}$ is set to half/single/double. $\kappa_{2}(\widetilde{A})$ is the same in all three cases. Note the different scales for the y axis.
@@
Table 3: LSIR iterations and the total count of FGMRES iterations within LSIR (Algorithm 4.1) for different values of $u_{s}$. Here $u$ is set to single, and $u_{s}$ is either half or single. - denotes that LSIR did not converge in 30 iterations. * denotes when $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES are set to double.
@@
Table 4: As in Table 3, but $u$ is set to double, and $u_{s}$ is either half, single, or double. - denotes that LSIR did not converge in 30 iterations. * denotes when $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES are set to quad and the maximum number of FGMRES iterations is increased to 80.